A Ba-135 atom (barium-135) has a nucleus with 56 protons and an atomic mass of 134.9056886 u. What is the nuclear binding energy per nucleon in this atom?

a) 1.22 MeV

b) 8.40 MeV

c) 14.4 MeV

d) 20.2 MeV

e) 1134 MeV
Answer #1

ne P2 = 56. n(n) = A-2 = 135-56 - 79, 5.68 1.007825= 56.4382 79x1.008665 = 79.684535 t. 136. 122735 - Mass def = Actual man m